OpenStudy (anonymous):

Multiply. state any excluded values. 5/2a . 3/a^2

OpenStudy (phi):

if this is the problem, \[ \frac{5}{2a} \cdot \frac{3}{a^2} \] multiply top times top and bottom times bottom because dividing by 0 is *very bad*, we do not allow a to be 0

OpenStudy (phi):

almost, but in the bottom you have a* a^2 remember that a^2 is short for a*a so this is a*a*a can you write this in short form ?

OpenStudy (phi):

\[ \frac{5}{2a} \cdot \frac{3}{a^2} = \frac{15}{2\cdot a\cdot a \cdot a} \]

OpenStudy (phi):

we do not want to divide by 0, so we do not want a=0, we say a≠0 that rules out choice c. the answer is \[ \frac{15}{2\cdot a\cdot a \cdot a} \] but you have to rewrite a*a*a in exponent form a times itself 3 times can be written how ? (other than a*a*a)

OpenStudy (phi):

why are you deleting your posts ?